OpenStudy (paperbacon):

Need help with algebra 2 workbook problems! Will fan and medal and testimonial!!!

OpenStudy (anonymous):

where are the problems

OpenStudy (paperbacon):

To 15

OpenStudy (paperbacon):

31-33

OpenStudy (paperbacon):

18 problems in all

OpenStudy (paperbacon):

@n648c788 can you help?

OpenStudy (paperbacon):

@Nnesha @mollyhatesmath

OpenStudy (anonymous):

yes I am reviewing some rules

OpenStudy (paperbacon):

Ok

OpenStudy (anonymous):

1.) amplitude is heights so it is 2 or -2 dosent matter period is 1/ frequency, so count wavelenghts 1/2=.5

OpenStudy (anonymous):

2.) is not periodic

OpenStudy (anonymous):

make sense

OpenStudy (paperbacon):

Yes, thank you!

OpenStudy (anonymous):

3-5 substact and add the value to 360 3-32+360 -32-360 make sense

OpenStudy (paperbacon):

What if its negative, then would I add?

OpenStudy (anonymous):

- - add numbers - + subtract numbers

OpenStudy (paperbacon):

Still confused. Could you give an example with end result so I can see the process?

OpenStudy (anonymous):

-32+360= 328 -32- 360=-396

OpenStudy (anonymous):

-229+360=131 -229-360=-589

OpenStudy (paperbacon):

Are thos both the answer. Like the answer is two numbers or pick the one thats positive?

OpenStudy (anonymous):

its both because it is not specify +, - value wanted

OpenStudy (anonymous):

5. 375 is not between 0 and 360

OpenStudy (paperbacon):

Ok, thanks! I think I got it now!

OpenStudy (anonymous):

Awseome!

OpenStudy (paperbacon):

Theres no answer for number then?

OpenStudy (paperbacon):

5

OpenStudy (anonymous):

yes good job

OpenStudy (paperbacon):

Thanks!

OpenStudy (paperbacon):

How do I do 6 through 8?

OpenStudy (anonymous):

6.) conversion degree to radians is angle *2Pi/360 -225*2pi/360= -3.92 radians terms of pie-1.25pi- we just did not multiply pie but we multiplied and divided

OpenStudy (anonymous):

7.) 120*2pi/360= 2.09 radians terms of pie= .667pi

OpenStudy (anonymous):

make sense

OpenStudy (paperbacon):

Yes, much!

OpenStudy (anonymous):

8.)600*2pi/360=10.47radians terms of pie= 3.33pi

OpenStudy (anonymous):

9-11 we are going backwards

OpenStudy (paperbacon):

Ok

OpenStudy (paperbacon):

Can you show me an example?

OpenStudy (anonymous):

9.) 5pi/6 *180/pi =150 degress 10.) -2.5pi *180/pi= -450 degress 11.) .8*pi/180 = .01396pi .01396pi*180/pi=5 degrees

OpenStudy (anonymous):

6-8 do value* pi/360

OpenStudy (paperbacon):

Okay, thanks! I understand how to do that from now on too!

OpenStudy (anonymous):

ok awseome!

OpenStudy (anonymous):

12. also?

OpenStudy (paperbacon):

I dont know about 12

OpenStudy (anonymous):

Ok because I can not see the graph

OpenStudy (anonymous):

13. then

OpenStudy (paperbacon):

Heres a better pic

OpenStudy (anonymous):

K

OpenStudy (anonymous):

1 cycle, amplitude is 2, period is 1

OpenStudy (anonymous):

make sense

OpenStudy (anonymous):

like when we did number 1

OpenStudy (paperbacon):

Yes

OpenStudy (anonymous):

ok next one

OpenStudy (paperbacon):

Ok

OpenStudy (anonymous):

y=4sin(2x) 4=amplitude sine period is 2Pi, period formula is 2pi/b, b=2 period=pi now graph

OpenStudy (anonymous):